OpenStudy (anonymous):

2√3x+2-1=5? can someone show me step by step how to do it?

Parth (parthkohli):

Is that the question?\[2\sqrt{3x + 2} - 1 = 5 \]

Parth (parthkohli):

I suppose that is.

OpenStudy (anonymous):

no there is more to it, i keep getting the wrong answer. The answer is 7/2 but i don't know how to get it.

Parth (parthkohli):

But is that the question? I can help you if that is.

OpenStudy (anonymous):

yes

Parth (parthkohli):

1) Add 1 to both sides. What do you get?

OpenStudy (anonymous):

6

Parth (parthkohli):

Yeah, right.\[ 2\sqrt{3x + 2} =6\]Now divide both sides by 2.

OpenStudy (anonymous):

ok √3x+2=3

OpenStudy (anonymous):

oh i see now, i missed a step. thanks for the help

Parth (parthkohli):

You're welcome!